A 46-centimeter-long board is cut into 3 pieces. The longest piece is twice as long as the shortest piece. The other piece is 6 cm longer that the shortest piece. How long is each piece?

Answer by checkley79(3341)

LET X EQUAL THE SHORTEST PIECE
THEN 2*X IS THE LONGEST PIECE
THEN X+6 IS THE THIRD PIECE
X+2X+(X+6)=46
4X+6=46
4X=46-6
4X=40
X=40/4
X=10 CM. FOR THE SHORTEST PIECE
2*10=20 CM. FOR THE LONGEST PIECE
10+6=16 CM. FOR THE THIRD PIECE
PROOF:
10+20+16=46
46=46